Portsmouth Company makes upholstered furniture. Its only variable cost is direct materials. The demand for the company's products far exceeds its manufacturing capacity. The bottleneck (or constraint) in the production process is upholstery labor-hours. Information concerning three of Portsmouth's upholstered chairs appears below:

Recliner Sofa Love Seat
Selling price per unit $ 1,362 $ 1,785 $ 1,625
Variable cost per unit $ 750 $ 1,200 $ 1,200
Upholstery labor-hours per unit 9 hours 13 hours 5 hours

Required:

1. Portsmouth is considering paying its upholstery laborers hourly compensation, in addition to their usual salaries, to work overtime. Assuming that this extra time would be used to produce sofas, up to how much of an overtime premium per hour should the company be willing to pay to keep the upholstery shop open after normal working hours?

2. A small nearby upholstering company has offered to upholster furniture for Portsmouth at a price of $40 per hour. The management of Portsmouth is confident that this upholstering company’s work is high quality and their craftsmen can work as quickly as Portsmouth’s own craftsmen on the simpler upholstering jobs such as the Love Seat. How much additional contribution margin per hour can Portsmouth earn if if it hires the nearby upholstering company to make Love Seats?

3. Should Portsmouth hire the nearby upholstering company?
Answer #1

SOLUTION

1.

Selling price per unit (A) 1,785
Less: Variable price per unit (B) 1,200
Contribution margin per unit (C=A-B) 585
Labor hours required for one unit (D) 13 hours
Contribution per labor hour (C/D) 45

2. Recliner

Selling price per unit (A) 1,362
Less: Variable price per unit (B) 750
Contribution margin per unit (C=A-B) 612
Labor hours required for one unit (D) 9 Hours
Contribution per labor hour (C/D) 68

Love seat

Selling price per unit (A) 1,625
Less: Variable price per unit (B) 1,200
Contribution margin per unit (C=A-B) 425
Labor hours required for one unit (D) 5 hours
Contribution per labor hour (C/D) 85
Price charged to upholster by another firm 40
Additional margin earned 45

3. Yes, it should hire the nearby upholstering company